Find vectors u,v is the element of R2 such that u is a scalar multiple of (1,3), v is orthogonal to (1,3), and (1,2) = u + v.


1
Expert's answer
2021-07-06T14:38:12-0400

Let "u=(u_1, u_2), v=(v_1, v_2)."

Given that "u" is a scalar multiple of "(1, 3)"


"u_1=k, u_2=3k"

Given that "v" is orthogonal  to "(1, 3)"


"v_1+3v_2=0"


Given that "u+v=(1, 2)"


"u_1+v_1=1""u_2+v_2=2"

We see that


"u_2=3u_1""v_1=-3v_2"

Substitute


"u_1-3v_2=1""3u_1+v_2=2"

Then


"v_2=-3u_1+2""u_1+9u_1-6=1"


"u_1=0.7""v_2=-0.1""u_2=2.1""v_1=0.3"

"u=(0.7, 2.1)"


"v=(0.3, -0.1)"